OpenStudy (anonymous):

I really need help please explain! Dustin is mixing concrete. A formula for concrete calls for 3/8 gal of water. Dustin wants to make 7/9 more concrete than the formula creates. How much water should Dustin use? Express your answer as a fraction in simplest form.

OpenStudy (anonymous):

How I see it is: 1 unit of concrete (or 9/9) calls for 3/8 gallon of water So to find out how much water you need for 1 and 7/9 units of concrete (or 16/9) Just multiply 3/8 by 16/9 = 48/72 (simplify) 2/3 gallon of water

OpenStudy (anonymous):

Um how did you get 16/9?

OpenStudy (anonymous):

1 unit of concrete = the fraction 9 over 9 (9/9) and you need any extra 7/9 unit of concrete So, i added 1 unit (9/9) to 7/9 of a unit of concrete to get 16/9 Lemme convert everything to decimals just to make sure it's right (everything is easier with decimals)

OpenStudy (anonymous):

Oh, ok now I get it

OpenStudy (anonymous):

Is that it or is there more to it?

OpenStudy (anonymous):

Ok, so 7/9 roughly gives you 0.78 + 1 = 1.78 (concrete) 3/8 = 0.375 (water for initial amount of concrete) 1.78 x 0.375 = 0.6675 (water for final amount of concrete) 2/3 (gallons of water calculated via fractions) = rough 0.667 So, I'm pretty certain 2/3 gallon of water is needed for an extra 7/9 more concrete

OpenStudy (anonymous):

Sum it all up, Dustin needs 2/3 gallon of water for the extra 7/9 concrete

OpenStudy (anonymous):

Thank you so much this really helps!